OpenStudy (anonymous):

directrix:y=-4,focus:(0,4)

Directrix (directrix):

 (x - 0)^2 = 16 (y - 0) x^2 = 16y y = (1/16) (x ^ 2) See link below for more about parabolas. http://www.purplemath.com/modules/parabola3.htm